# Limits of Probability Measures with General Coefficients

### v2

@@ 3.5. Sequences involving changes
It suffices to show that if $s\in T^{\prime}$, for $d>\lambda_{1}+\cdots+\lambda_{j}$, $s_{d}$ is not a change from $i_{d}$ to $x_{i}$ or a term multiplication that degree-alters $x_{i}$, $N-j+1\leq i\leq N$. For the sake of contradiction, assume $s\in T^{\prime}$ and $s_{d}$ contains $x_{i}$, $N-j+1\leq i\leq N$, for some $d>\lambda_{1}+\cdots+\lambda_{j}$. From the conditions of Lemma 3.16, $x_{i}$ must be degree-altered by a term multiplication. After, $x_{i}$ will be in all terms of $r(s)_{d}$; note that in the term multiplication, we do not take the derivative with respect to $x_{i}$ because $d>\lambda_{1}+\cdots+\lambda_{j}$. But, $x_{i}$ cannot be removed from a term by an operator unless the operator converts the term to zero, a contradiction to $s\in T_{N,\theta_{N}}^{2}(\mathfrak{r})$. ∎
@@
Note that because $T^{\prime}\subset T\cap T_{N,\theta_{N}}^{2}(\mathfrak{r})$, $U^{\prime}\subset U\cap T$. Also, $U\cap T\backslash U^{\prime}$ is the set of $s\in T\cap T^{2}_{N,\theta_{N}}\backslash T^{\prime}$ such that $r(s)_{k}=dp$, $d\not=0$. However, the number of $s\in U\backslash U^{\prime}$ with $l$ changes is $O(N^{l-1})$ from the proof of Lemma 3.13 without derivatives and using changes instead of switches, as in the proof of Lemma 3.16. Therefore, the number $s\in U\cap T\backslash U^{\prime}$ with $l$ changes is $O(N^{l-1})$. Then, by following the previous argument, we get that the coefficient of $p$ from sequences $s\in T\cap T_{N,\theta_{N}}^{2}\backslash T^{\prime}$ is $O(N^{(k-m)(1-c)-1})$. Thus,
@@
3.6. Proof of Theorem 3.2
Suppose $\mathfrak{r}$ is the sequence that is described in the statement of Proposition 3.12. The result implies that as $N\rightarrow\infty$,
@@
using (16) with $k=|\lambda|(1-c)-1$ and $\epsilon=1$. Afterwards, from Theorem 3.17, Theorem 3.20, and (16),
@@
Recall the definition of $\vec{X}_{i}$ from the statement of Theorem 3.20. To evaluate this, note that because $F_{N}$ is symmetric,
@@
and the limiting sequence of (18) with respect to $1$ for $1\leq i\leq m$ is $f=\{c^{d,\nu}\}_{d\geq 0,\nu\in P}$, where $c^{d,\nu}=(d+1)c_{\nu+(d+1)}$. From Theorem 2.13 with $\lim_{N\rightarrow\infty}N^{c}\theta_{N}=\theta$,
@@ 3.7. Example
for all $\nu\in P^{+}$. From Theorem 3.2, for $\lambda\in P^{+}$,
@@ 4.1. Bessel generating functions
For $\theta\in\mathbb{C}$ and $a_{1},\ldots,a_{N},x_{1},\ldots,x_{N}\in\mathbb{C}$ such that the $a_{i}$ are not all zero, the function $\text{Exp}_{(a_{1},\ldots,a_{N})}(x_{1},\ldots,x_{N};\theta):\mathbb{C}^{N}\backslash\{0\}\times\mathbb{C}^{N}\times\mathbb{C}\rightarrow\mathbb{C}$ is the unique meromorphic function such that for all $(a_{1},\ldots,a_{N})\in\mathbb{C}^{N}\backslash\{0\}$ and $\theta\in\mathbb{C}$, $\mathcal{D}_{i}^{\theta}\text{Exp}_{(a_{1},\ldots,a_{N})}(\cdot;\theta)=a_{i}\text{Exp}_{(a_{1},\ldots,a_{N})}(\cdot;\theta)$ for $1\leq i\leq N$ and $\text{Exp}_{(a_{1},\ldots,a_{N})}(0;\theta)=1$. The multivariate Bessel function $B_{(a_{1},\ldots,a_{N})}(x_{1},\ldots,x_{N};\theta):\mathbb{C}^{N}\backslash\{0\}\times\mathbb{C}^{N}\times\mathbb{C}\rightarrow\mathbb{C}$ is defined as
$B_{(a_{1},\ldots,a_{N})}(x_{1},\ldots,x_{N};\theta)\triangleq\frac{1}{N!}\sum_{\sigma\in S_{N}}\text{Exp}_{(a_{1},\ldots,a_{N})}(x_{\sigma(1)},\ldots,x_{\sigma(N)};\theta),$
where $S_{N}$ is the set of permutations of $[N]$.
@@
The following result is well-known and is true for all $\theta\in\mathbb{C}$; it is straightforward to prove using Definition 4.1.
@@
The Bessel generating function is also studied in [betaprocess, airy, matrix, rectangularmatrix]. A particular example that we consider is the Bessel generating function of the $\beta$-Hermite ensemble, see Sections 5 and 6.
From Lemma 1.5, if $\mu\in\mathcal{M}_{N}$ is exponentially decaying, then the Bessel generating function of $\mu$ converges in a neighborhood of the origin in $\mathbb{C}^{N}$. Additionally, from Proposition 4.2,
@@
Note that the paper [matrix] proves Lemma 1.5 and Proposition 4.4 for when $\theta>0$ using the integral formulation of the multivariate Bessel function as the expectation of a function over a probability distribution over the Gelfand-Tsetlin patterns. For the regime $(\theta)\geq 0$, we can prove the results using the same method as the paper as well as the inequality
$|\text{Exp}_{(a_{1},\ldots,a_{N})}(x_{1},\ldots,x_{N};\theta)|\leq\sqrt{N!}\exp\left(\max_{\sigma\in S_{N}}\left(\sum_{i=1}^{N}(a_{i}x_{\sigma(i)})\right)\right)$
for all $a_{1},\ldots,a_{N},x_{1},\ldots,x_{N}\in\mathbb{C}$, see [dunklbound].
@@
From Proposition 4.4, for exponentially decaying $\mu\in\mathcal{M}_{N}$ and $\lambda\in P^{+}$,
@@
A step for showing that $\{\mu_{N}\}_{N\geq 1}$ satisfy a $c$-LLN is to use (19) with $\mu=\mu_{N}$ and $\theta=\theta_{N}$, see Section 4.2.
@@
Next, suppose $\lambda\in P^{+}$ has length $m$. Suppose $\mathfrak{r}=\{i_{j}\}_{1\leq j\leq|\lambda|}$ is a sequence where the first $\lambda_{1}$ values are $1$, the next $\lambda_{2}$ are $2$, and so forth, until the last $\lambda_{m}$ are $m$. For Theorem 1.7, we look at the limit of (19) for $\theta=\theta_{N}$ and $\mu=\mu_{N}$ as $N\rightarrow\infty$. For $N\geq 1$, we let $F_{N}(x_{1},\ldots,x_{N})$ be a function such that $\exp(F_{N}(x_{1},\ldots,x_{N}))=G_{\theta_{N}}(x_{1},\ldots,x_{N};\mu_{N})$; because the $\mu_{N}$ is exponentially decaying, by Lemma 1.5, $G_{\theta_{N}}(\cdot;\mu_{N})$ is holomorphic in a neighborhood of the origin and furthermore evaluates to one at the origin, and is therefore nonzero in a neighborhood of the origin as well. Hence, there exists $F_{N}$ that is also symmetric and holomorphic in a neighborhood of the origin. Furthermore, we can assume that $F_{N}(0,\ldots,0)$ equals $\ln(G_{\theta_{N}}(0,\ldots,0;\mu_{N}))=0$, rather than a multiple of $2\pi i$.
Suppose that $\tilde{F}_{N}(x_{1},\ldots,x_{N})$ is the polynomial for $F_{N}(x_{1},\ldots,x_{N})$ from Proposition 4.6 with $k=|\lambda|$. By Proposition 4.6, (19), and Lemma 4.7,
@@
The symmetric polynomial $\tilde{F}_{N}(x_{1},\ldots,x_{N})$ is a symmetric formal series for $N\geq 1$, and with the Theorem 1.7 conditions, Theorem 3.2 can be used on $\{\tilde{F}_{N}(x_{1},\ldots,x_{N})\}_{N\geq 1}$. This completes the proof.
@@ 5. Eigenvalue Distributions
(20) $d_{N,\beta}(x_{1},\ldots,x_{N})=C_{N,\beta}\prod_{1\leq i<j\leq N}|x_{i}-x_{j}|^{\beta}\prod_{i=1}^{N}e^{-\frac{\beta x_{i}^{2}}{4}},$
where
@@
Suppose $\{\mu_{N}\}_{N\geq 1}$ is a sequence of probability measures such that $\mu_{N}$ is in $\mathcal{M}_{N}$ and has density $d^{s}_{N,\beta}$. After scaling by $N^{-k}$, the $2k$th moment of the $\beta$-Hermite ensemble converges to the $k$th Catalan number for $k\geq 0$, see [eigenvalstat]*Theorem 6.2.3. Proposition 5.3 shows that this is the case and proves 00-LLN satisfaction as an example of an application of Theorem 1.7.
@@
Let $a_{1},\ldots,a_{N},y_{1},\ldots,y_{N},c$ be $12N+$ 1 arbitrary complex numbers. Then, $B_{(ca_{1},\ldots,ca_{N})}(y_{1},\ldots,y_{N})=B_{(a_{1},\ldots,a_{N})}(cy_{1},\ldots,cy_{N})$.
@@
With Lemma 5.1,
[…]
@@
Suppose $\theta>0$. Consider the sequence $\{\mu_{N}\}_{N\geq 1}$ of probability measures such that for all positive integers $N$, $\mu_{N}\in\mathcal{M}_{N}$ and $\mu_{N}$ has probability density $d_{N,2\theta}^{s}$ from (21). Then, $\{\mu_{N}\}_{N\geq 1}$ satisfy a 00-LLN, with $m_{2k-1}=0$ and $m_{2k}=\frac{1}{k+1}\binom{2k}{k}$ for $k\geq 1$.
@@
Then, from Lemma 5.2 and the previous equation,
@@
Afterwards, by Theorem 1.7 with $c=0$ and $\theta_{N}=\theta$ for $N\geq 1$, $\{\mu_{N}\}_{N\geq 1}$ satisfy a LLN with $c_{(2)}=\frac{1}{2\theta}$ and $c_{\nu}=0$ for $\nu\in P^{+}$ such that $\nu\not=(2)$. For $k\geq 1$, $m_{2k-1}=0$ and $m_{2k}$ is the number of $\pi\in NC(2k)$ which have all blocks of size $2$, which is the $k$th Catalan number and equals $\frac{1}{k+1}\binom{2k}{k}$. ∎
From Proposition 5.3, if $(a_{1},\ldots,a_{N})$ is distributed with density $d_{N,\beta}(x_{1},\ldots,x_{N})$, for a positive integer $s$ and positive integers $k_{i}$, $1\leq i\leq s$,
@@
with the moments $m_{k}$ given in Proposition 5.3. Thus, the distributions $d_{N,\beta}$ satisfy a $\frac{1}{2}$-LLN, although note that $\theta$ is constant rather than proportional to $N^{-\frac{1}{2}}$. Sometimes, as seen above, for sequences $\{\mu_{N}\}_{N\geq 1}$ of exponentially decaying probability measures, we can scale $\mu_{N}$ by an appropriate power of $N$ for $N\geq 1$ and use Lemma 5.2 to satisfy the conditions of Theorem 1.7. We state this idea rigorously in the following result.
@@
For $D=N^{\Delta}$, let $\mu_{N}^{D}$ be the pushforward of $\mu_{N}$ with respect to the function $f:A\rightarrow A,x\mapsto Dx$. By Lemma 5.2, for all $\nu\in P^{+}$,
@@
for all positive integers $i_{1},\ldots,i_{r}$ such that $\sigma((i_{1},\ldots,i_{r}))=\nu$. Then, from Theorem 1.7, $\{\mu_{N}^{D}\}_{N\geq 1}$ satisfy a $c$-LLN with the moments $\{m_{k}\}_{k\geq 1}$. Then, $\{\mu_{N}\}_{N\geq 1}$ satisfy a $(c+\Delta)$-LLN with these moments. ∎

### v3

@@ 3.5. Sequences involving changes
It suffices to show that if $s\in T^{\prime}$, for $d>\lambda_{1}+\cdots+\lambda_{j}$, $s_{d}$ is not a change from $i_{d}$ to $x_{i}$ or a term multiplication that degree-alters $x_{i}$, $N-j+1\leq i\leq N$. For the sake of contradiction, assume $s\in T^{\prime}$ and $s_{d}$ contains $x_{i}$, $N-j+1\leq i\leq N$, for some $d>\lambda_{1}+\cdots+\lambda_{j}$. From the conditions of Lemma˜3.16, $x_{i}$ must be degree-altered by a term multiplication. After, $x_{i}$ will be in all terms of $r(s)_{d}$; note that in the term multiplication, we do not take the derivative with respect to $x_{i}$ because $d>\lambda_{1}+\cdots+\lambda_{j}$. But, $x_{i}$ cannot be removed from a term by an operator unless the operator converts the term to zero, a contradiction to $s\in T_{N,\theta_{N}}^{2}(\mathfrak{r})$. ∎
@@
[…]
@@
3.6. Proof of Theorem˜3.2
Suppose $\mathfrak{r}$ is the sequence that is described in the statement of Proposition˜3.12. The result implies that as $N\rightarrow\infty$,
@@
using (15) with $k=|\lambda|(1-c)-1$ and $\epsilon=1$. Afterwards, from Theorem˜3.17, Theorem˜3.20, and (15),
@@
Recall the definition of $\vec{X}_{i}$ from the statement of Theorem˜3.20. To evaluate this, note that because $F_{N}$ is symmetric,
@@
and the limiting sequence of (17) with respect to $1$ for $1\leq i\leq m$ is $f=\{c^{d,\nu}\}_{d\geq 0,\nu\in P}$, where $c^{d,\nu}=(d+1)c_{\nu+(d+1)}$. From Theorem˜2.13 with $\lim_{N\rightarrow\infty}N^{c}\theta_{N}=\theta$,
@@ 3.7. Example
for all $\nu\in P^{+}$. From Theorem˜3.2, for $\lambda\in P^{+}$,
@@ 4.1. Bessel generating functions
For $\theta\in\mathbb{C}$ and $a_{1},\ldots,a_{N},x_{1},\ldots,x_{N}\in\mathbb{C}$ such that the $a_{i}$ are not all zero, the function $\text{Exp}_{(a_{1},\ldots,a_{N})}(x_{1},\ldots,x_{N};\theta):\mathbb{C}^{N}\times\mathbb{C}^{N}\times\mathbb{C}\rightarrow\mathbb{C}$ is the unique meromorphic function such that for all $(a_{1},\ldots,a_{N})\in\mathbb{C}^{N}$ and $\theta\in\mathbb{C}$, $\mathcal{D}_{i}^{\theta}\text{Exp}_{(a_{1},\ldots,a_{N})}(\cdot;\theta)=a_{i}\text{Exp}_{(a_{1},\ldots,a_{N})}(\cdot;\theta)$ for $1\leq i\leq N$ and $\text{Exp}_{(a_{1},\ldots,a_{N})}(0;\theta)=1$. The multivariate Bessel function $B_{(a_{1},\ldots,a_{N})}(x_{1},\ldots,x_{N};\theta):\mathbb{C}^{N}\times\mathbb{C}^{N}\times\mathbb{C}\rightarrow\mathbb{C}$ is defined as
$B_{(a_{1},\ldots,a_{N})}(x_{1},\ldots,x_{N};\theta)\triangleq\frac{1}{N!}\sum_{\sigma\in\mathcal{S}_{N}}\text{Exp}_{(a_{1},\ldots,a_{N})}(x_{\sigma(1)},\ldots,x_{\sigma(N)};\theta),$
where $\mathcal{S}_{N}$ is the set of permutations of $[N]$.
@@
The following result is well-known and is true for all $\theta\in\mathbb{C}$; it is straightforward to prove using Definition˜4.1.
@@
Remark 4.3.
It is not immediate that the multivariate Bessel function is the only solution of the system of partial differential equations in Proposition˜4.2. This is established when $\theta$ has certain properties in [dunklbessel]*Section 5.
@@
The Bessel generating function is also studied in [dunklbound, betaprocess, gaussianfluctuation, airy, matrix, rectangularmatrix]. A particular example that we consider is the Bessel generating function of the $\beta$-Hermite ensemble, see Sections˜5 and 6.
From Lemma˜1.4, if $\mu\in\mathcal{M}_{N}$ is exponentially decaying, then the Bessel generating function of $\mu$ converges in a neighborhood of the origin in $\mathbb{C}^{N}$. Additionally, from Proposition˜4.2,
@@
Note that the paper [matrix] proves Lemma˜1.4 and Proposition˜4.5 for when $\theta>0$ using the integral formulation of the multivariate Bessel function as the expectation of a function over a probability distribution over the Gelfand-Tsetlin patterns. For the regime $(\theta)\geq 0$, we can prove the results using the same method as the paper as well as the inequality
(18) $\begin{split}|\text{Exp}_{(a_{1},\ldots,a_{N})}(x_{1},\ldots,x_{N};\theta)|&\leq\sqrt{N!}\exp\left(\max_{\sigma\in\mathcal{S}_{N}}\left(\sum_{i=1}^{N}(a_{i}x_{\sigma(i)})\right)\right)\\ &\leq\sqrt{N!}\exp(|(a_{1},\ldots,a_{N})|\cdot|(x_{1},\ldots,x_{N})|)\end{split}$
for all $a_{1},\ldots,a_{N},x_{1},\ldots,x_{N}\in\mathbb{C}$ from [dunklbound]*Corollary 3.2. The required bounds on the derivatives of the multivariate Bessel function for the proof of Proposition˜4.5 follow from [dunklbound]*Lemma 3.5.
@@
From Proposition˜4.5, for exponentially decaying $\mu\in\mathcal{M}_{N}$ and $\lambda\in P^{+}$,
@@
A step for showing that $\{\mu_{N}\}_{N\geq 1}$ satisfy a $c$-LLN is to use (19) with $\mu=\mu_{N}$ and $\theta=\theta_{N}$, see Section˜4.2.
@@
Next, suppose $\lambda\in P^{+}$ has length $m$. Suppose $\mathfrak{r}=\{i_{j}\}_{1\leq j\leq|\lambda|}$ is a sequence where the first $\lambda_{1}$ values are $1$, the next $\lambda_{2}$ are $2$, and so forth, until the last $\lambda_{m}$ are $m$. For Theorem˜1.6, we look at the limit of (19) for $\theta=\theta_{N}$ and $\mu=\mu_{N}$ as $N\rightarrow\infty$. For $N\geq 1$, we let $F_{N}(x_{1},\ldots,x_{N})$ be a function such that $\exp(F_{N}(x_{1},\ldots,x_{N}))=G_{\theta_{N}}(x_{1},\ldots,x_{N};\mu_{N})$; because the $\mu_{N}$ is exponentially decaying, by Lemma˜1.4, $G_{\theta_{N}}(\cdot;\mu_{N})$ is holomorphic in a neighborhood of the origin and furthermore evaluates to one at the origin, and is therefore nonzero in a neighborhood of the origin as well. Hence, there exists $F_{N}$ that is also symmetric and holomorphic in a neighborhood of the origin. Furthermore, we can assume that $F_{N}(0,\ldots,0)$ equals $\ln(G_{\theta_{N}}(0,\ldots,0;\mu_{N}))=0$, rather than a multiple of $2\pi i$.
Suppose that $\tilde{F}_{N}(x_{1},\ldots,x_{N})$ is the polynomial for $F_{N}(x_{1},\ldots,x_{N})$ from Proposition˜4.7 with $k=|\lambda|$. By Proposition˜4.7, (19), and Lemma˜4.8,
@@
The symmetric polynomial $\tilde{F}_{N}(x_{1},\ldots,x_{N})$ is a symmetric formal series for $N\geq 1$, and with the Theorem˜1.6 conditions, Theorem˜3.2 can be used on $\{\tilde{F}_{N}(x_{1},\ldots,x_{N})\}_{N\geq 1}$. This completes the proof.
@@ 5. Eigenvalue Distributions
(20) $d_{N,\beta}(x_{1},\ldots,x_{N})=C_{N,\beta}\prod_{1\leq i<j\leq N}|x_{i}-x_{j}|^{\beta}\prod_{i=1}^{N}e^{-\frac{\beta x_{i}^{2}}{4}}$
over $\mathbb{R}^{n}$, where
@@
Suppose $\{\mu_{N}\}_{N\geq 1}$ is a sequence of probability measures such that $\mu_{N}$ is in $\mathcal{M}_{N}$ and has density $d^{s}_{N,\beta}$. After scaling by $N^{-k}$, the $2k$th moment of the $\beta$-Hermite ensemble converges to the $k$th Catalan number for $k\geq 0$, see [eigenvalstat]*Theorem 6.2.3. Proposition˜5.3 shows that this is the case and proves 0-LLN satisfaction as an example of an application of Theorem˜1.6.
@@
Let $a_{1},\ldots,a_{N},y_{1},\ldots,y_{N},c$ be $2N+1$ arbitrary complex numbers and suppose $\theta>0$. Then, $B_{(ca_{1},\ldots,ca_{N})}(y_{1},\ldots,y_{N};\theta)=B_{(a_{1},\ldots,a_{N})}(cy_{1},\\ \ldots,cy_{N};\theta)$.
@@
With Lemma˜5.1,
$G_{\theta}(x_{1},\ldots,x_{N};\mu_{c})$ $=\int_{a\in A}B_{a}(x_{1},\ldots,x_{N};\theta)d\mu_{c}(a_{1},\ldots,a_{N})$
$=\int_{a^{\prime}\in A}B_{ca^{\prime}}(x_{1},\ldots,x_{N};\theta)d\mu(a_{1}^{\prime},\ldots,a_{N}^{\prime})$
$=\int_{a^{\prime}\in A}B_{a^{\prime}}(cx_{1},\ldots,cx_{N};\theta)d\mu(a_{1}^{\prime},\ldots,a_{N}^{\prime})$
@@
Suppose $\theta>0$. Consider the sequence $\{\mu_{N}\}_{N\geq 1}$ of probability measures such that for all positive integers $N$, $\mu_{N}\in\mathcal{M}_{N}$ and $\mu_{N}$ has probability density $d_{N,2\theta}^{s}$ from (21). Then, $\{\mu_{N}\}_{N\geq 1}$ satisfy a 0-LLN, with $m_{2k-1}=0$ and $m_{2k}=\frac{1}{k+1}\binom{2k}{k}$ for $k\geq 1$.
@@
Then, from Lemma˜5.2 and the previous equation,
@@
Afterwards, by Theorem˜1.6 with $c=0$ and $\theta_{N}=\theta$ for $N\geq 1$, $\{\mu_{N}\}_{N\geq 1}$ satisfy a 0-LLN with $c_{(2)}=\frac{1}{2\theta}$ and $c_{\nu}=0$ for $\nu\in P^{+}$ such that $\nu\not=(2)$. For $k\geq 1$, $m_{2k-1}=0$ and $m_{2k}$ is the number of $\pi\in NC(2k)$ which have all blocks of size $2$, which is the $k$th Catalan number and equals $\frac{1}{k+1}\binom{2k}{k}$. ∎
From Proposition˜5.3, if $(a_{1},\ldots,a_{N})$ is distributed with density $d_{N,\beta}(x_{1},\ldots,x_{N})$, for a positive integer $s$ and positive integers $k_{i}$, $1\leq i\leq s$,
@@
with the moments $m_{k}$ given in Proposition˜5.3. Thus, the distributions $d_{N,\beta}$ satisfy a $\frac{1}{2}$-LLN, although note that $\theta$ is constant rather than proportional to $N^{-\frac{1}{2}}$. Sometimes, as seen above, for sequences $\{\mu_{N}\}_{N\geq 1}$ of exponentially decaying probability measures, we can scale $\mu_{N}$ by an appropriate power of $N$ for $N\geq 1$ and use Lemma˜5.2 to satisfy the conditions of Theorem˜1.6. We state this idea rigorously in the following result.
@@
For $D=N^{\Delta}$, let $\mu_{N}^{D}$ be the pushforward of $\mu_{N}$ with respect to the function $f:A\rightarrow A,x\mapsto Dx$. By Lemma˜5.2, for all $\nu\in P^{+}$,
@@
for all positive integers $i_{1},\ldots,i_{r}$ such that $\sigma((i_{1},\ldots,i_{r}))=\nu$. Then, from Theorem˜1.6, $\{\mu_{N}^{D}\}_{N\geq 1}$ satisfy a $c$-LLN with the moments $\{m_{k}\}_{k\geq 1}$. Then, $\{\mu_{N}\}_{N\geq 1}$ satisfy a $(c+\Delta)$-LLN with these moments. ∎